Two vessel of capacity 1.5 ltr and 2ltr contain hydrogen gas and oxygen gas respectively under a pressure of 750mm and 100mm. The gases are mixed together in a 5ltr vessel.What will be the final pressure of mixture?

Grade:11

1 Answers

Vikas TU
7 years ago
Remaining parameters that is Temperature, moles to be constant.
From Boyle’s law,
P1V1 + P2V2 = PV
Hence,
750*1.5 + 100*2 = P*5
200 + 1125 = 5P
P = 1325/5 = 265 mm.